the difference of two numbers is 5. when twice the larger is subtracted from 6 times the smaller, the difference is 26. find each number.
Answer by mananth(16946)

1 x + -1 y = 5 .............1
-2 x + 6 y = 26 .............2
Eliminate y
multiply (1)by 6
Multiply (2) by 1
6 x -6 y = 30
-2 x + 6 y = 26
Add the two equations
4 x = 56
/ 4
x=14
plug value of x in (1)
x -1 y = 5
14 -1 y = 5
-1 y = 5-14
-1 y = -9
y = 9